# Deep learning inference of the neutron star equation of state

### v1

@@ III.3 Bayesian deep network
IV Conclusions
@@
There are various interesting future directions for our work. First, one could allow the observational data to have a lower or varying uncertainties. We expect that such step would allow for a higher accuracy of our network. It would also be interesting to perform a similar analysis employing the machinery of normalising flows to construct the posterior of our data, thus providing a conceptually different analysis of the pipeline presented here. Our pipeline was based on a particular parametrisation for the EOS, and it would be essential to test the performance of our network on different EOS parametrisations for the high-density regime of the star. Furthermore, to reduce the bias in our training data toward lower values of the vacuum energy, it would be crucial to allow for more configurations with positive values of vacuum energy shift. Finally, a promising direction would be to complement our datasets with gravitational waveforms, this way extending our analysis to the case of neutron-star mergers. We leave these aspects for future work.
Acknowledgement
We thank the EuCAIFCon for useful discussion on this work during the conference. G.V. acknowledges support by the Czech Academy of Sciences under the grant number LQ100102101, and I.D.S by the Czech Grant Agency (GAĈR) under the grant number 21-16583M.

### v2

@@ 3.3 Bayesian deep network
4 Employing real observational data
We now apply our networks to analyze real observational data. Since we are currently lacking varied and precise observations on the tidal deformability [^5] The current constraint on the tidal deformability from GW170817 for a neutron star with $1.4\,M_{\odot}$ are $\lambda(1.4\,M_{\odot})\leq 800$, assuming a low-spin prior [19]., we restrict to the analysis of masses and radii only. We use measures from X-ray observations, and we consider two different sources according to their catalogued names:
  • 6 thermonuclear bursters: 4U 1608-52, 4U 1724-207, KS 1731-260, EXO 1745-248, SAX J1748.9-2021 and 4U 1820-30 [7, 9].
  • 2 pulsars from NICER: PSR J0030+0451 [64] and PSR J0740+6620 [18].
From these observations we randomly sample 15 $(M,R)$ points and we appropriately normalized them by $M_{\text{norm}}=3\,M_{\odot}$ and $R_{\text{norm}}=20\,\text{km}$, as described in Section 3.1. We used these data as input for both the deterministic regression model and the Bayesian network, discussed in Sections 3.2 and 3.3. We show the results in Fig. 12. On the left panel, we show a statistical sample of the predictions as outputted from the Bayesian deep network for $n_{\text{samples}}=100$ realisations of the output (coloured points). The black dots and the associated error bars are computed as the average and $2\sigma$ error bars respectively, given the statistical sample for each column of the output’s array. We also show the speed-of-sound profile as outputted from the deterministic model (red dashed line). We can see that the results from the deterministic network are in agreement within $2\,\sigma$ with those from the probabilistic one. On the right panel, we show the predictions for the vacuum energy associated with the phase transition in the stellar core. The histograms are based on the predictions of the Bayesian deep network, for $n_{\text{samples}}=100$ realisations of the output. We also show the mean value with a $1\sigma$ interval predicted from the probabilistic network, i.e. $\text{sign}(\Lambda)\cdot|\Lambda|^{1/4}=-61\pm 29\,\text{MeV}$, represented by a black solid and dashed line respectively. We include the result from the deterministic network, i.e. $\Lambda=-(75\,\text{MeV})^{4}$, represented by a red line. We can see that the predictions from the deterministic network are in agreement with those from the probabilistic one. However, we refrain from making any claims on the value of $\Lambda$, since as we saw during the training and testing of our models, the networks cannot confidently constrain this value. Improving both the training dataset and the network can potentially shed some light on possible constraints on the value of $\Lambda$. We leave this for future work. Finally, we also note that the approximate $2\,\sigma$ value of $\Lambda$ is practically compatible with $\Lambda=0$.
Figure 12: Left panel: statistical sample of the predictions as outputted from the Bayesian deep network for $n_{\text{samples}}=100$ realisations of the output (coloured points). The black dots and the associated error bars are computed as the average and $2\sigma$ error bars respectively, given the statistical sample for each column of the output’s array. We also show the speed-of-sound profile as outputted from the deterministic model (red dashed line). We can see that the results from the deterministic network are in agreement within $2\,\sigma$ with those from the probabilistic one. We note that both the Bayesian and the deterministic deep networks predict a maximum value of speed-of-sound $c_{s}\simeq 0.8$ at $\rho/\rho_{f}\simeq 0.2$. Such large value of speed-of-sound seems to be less favourable from previous results in the literature, see e.g. Ref. [74] where they compute the probability density function of $c_{s}^{2}$. We found that our networks’ outputs consistently show a bias towards large maximum values of $c_{s}$ at $\rho/\rho_{f}\simeq 0.2$ during the training process for all cases considered. Of course, the large value at $\rho/\rho_{f}\simeq 0.2$ refers to the computed mean of the sound speed, and the corresponding error bar needs be accounted for too. This large mean value appears nevertheless to be somewhat intriguing and deserves a further investigation in the future. Right panel: predictions for the vacuum energy associated with the phase transition in the stellar core. The histograms are based on the predictions of the Bayesian deep network, for $n_{\text{samples}}=100$ realisations of the output. We also show the mean value with a $1\,\sigma$ interval predicted from the probabilistic network, i.e. $\text{sign}(\Lambda)\cdot|\Lambda|^{1/4}=-61\pm 29\,\text{MeV}$, represented by a black solid and dashed line respectively. We include the result from the deterministic network, i.e. $\Lambda=-(75\,\text{MeV})^{4}$, represented by a red line. The predictions from the deterministic network are in agreement with those from the probabilistic one. Notice that the approximate $2\sigma$ value of $\Lambda$ is practically compatible with $\Lambda=0$. However, given the simplifying assumptions of our network and the small number of real observations, these results need to be treated with extreme care.
5 Conclusions
@@
Finally, we applied our network to real mass and radius data from X-ray observations. We used these data as input to both our deterministic and probabilistic network. The results are shown in Fig 12, where we show both the output for the speed-of-sound profile and that for the vacuum energy shift. The results from the deterministic network and those from the probabilistic one are in agreement. From the Bayesian network, we find that the vacuum energy jump’s mean value with a $1\sigma$ interval is given by $\text{sign}(\Lambda)\cdot|\Lambda|^{1/4}=-61\pm 29\,\text{MeV}$. Given the simplifying assumptions of our network and the small number of real observations, this result needs to be treated with extreme care. We also notice that the approximate $2\,\sigma$ value of $\Lambda$ is practically compatible with $\Lambda=0$.”
There are various interesting future directions for our work. First, one could allow the observational data to have a lower or varying uncertainties. We expect that such step would allow for a higher accuracy of our network. It would also be interesting to perform a similar analysis employing the machinery of normalising flows to construct the posterior of our data, thus providing a conceptually different analysis of the pipeline presented here. Our pipeline was based on a particular parametrisation for the EOS, and it would be essential to test the performance of our network on different EOS parametrisations for the high-density regime of the star. We also intend to improve the study at low-densities by either allowing for a wider range of low-density EOSs from those commonly used in the literature, or by extending the analysis of the regression model to take into account this region as well. This would allow for a more appropriate treatment of the uncertainties that are still present at low-densities. Furthermore, to reduce the bias in our training data toward lower values of the vacuum energy, it would be crucial to allow for more configurations with positive values of vacuum energy shift. Finally, a promising direction would be to complement our datasets with gravitational waveforms, this way extending our analysis to the case of neutron-star mergers. We leave these aspects for future work.
Acknowledgments
We thank the referee for their constructive feedback. We thank the EuCAIFCon for useful discussion on this work during the conference. G.V. acknowledges support by the Czech Academy of Sciences under the grant number LQ100102101, and I.D.S by the Czech Grant Agency (GAĈR) under the grant number 21-16583M.